OpenStudy (anonymous):

Create your own word problem that can be solved using a system of equations. It should not be a problem used in the notes or activities in this module. You will need to type the problem, the system, and the solution to receive full credit.

OpenStudy (anonymous):

Please help!

OpenStudy (puzzler7):

Megan has 5 apples and bananas altogether. She has one more banana than apple. System:\[a+b=5\]\[b=a+1\]Solution: \[a+(a+1)=5\]\[2a+1=5\]\[2a=4\]\[a=2\]\[2+b=5\]\[b=3\]

OpenStudy (puzzler7):

a=2, b=3

OpenStudy (anonymous):

thank you so much for your help!!

OpenStudy (puzzler7):

No need to thsnk me. Just medal/fan.